OpenStudy (anonymous):

simple math question: in a substitution (calculus) where du=(x-1)dx, can I add 1 one both sides to get du+1=xdx ? or do the dx prevent it since it multiplies (x-1) ?

OpenStudy (lgbasallote):

if you expand du = (x - 1)dx, you should get du = xdx - dx (according to distributive property) does that answer your question?

OpenStudy (anonymous):

yes thank you !